OpenStudy (anonymous):

What is the approximate radius of a sphere whose volume is 1,349 cm³?

OpenStudy (anonymous):

5.7 cm 6.9 cm 11 cm 14.7 cm one of theses answers is correct.

OpenStudy (anonymous):

volume of sphere = (4/3)pi r^3 = 1349 r^3 = 1349 / (4pi/3) = 3*1349 ------ 4pi r = cuberoot of (3*1349 / 4pi) = cuberrot of 322 = 6.9

OpenStudy (anonymous):

Thank you soooo much!!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i like ur name...